OpenStudy (mrs.ambrose614):

Question 11 (Multiple Choice Worth 1 points) (03.03 MC) The equation below represents Function A and the graph represents Function B: Function A f(x) = – 2x + 1 Function B graph of line going through ordered pairs negative 1, negative 5 and 2, 1 and 3, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 x Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 x Slope of Function B Slope of Function B = – Slope of Function A @will.h

OpenStudy (will.h):

Function A : f(x) = -2x + 1 clearly the slope is -2 Function B (-1,-5)(2,1)(3,3) we could pick any 2 points of those to find the slope slope = y2-y1/x2-x1 slope = 3 - 1/3-2 slope = 2

OpenStudy (mrs.ambrose614):

So the awnser would be b or c???????

OpenStudy (will.h):

neither could you double check what option D is?

OpenStudy (will.h):

cuz it doesn't make sense

OpenStudy (mrs.ambrose614):

D is slope of function b equals -slope of function a

OpenStudy (will.h):

and that's indeed the case

OpenStudy (mrs.ambrose614):

So d would be correct???

OpenStudy (will.h):

yes because slope of B is 2 and slope of A is -2 thus slope of B = negative the slope of A see if that's correct -(-2) = 2 thus yeah it is true

OpenStudy (mrs.ambrose614):

Thanks so much smarty pants lol @Will.H

OpenStudy (will.h):

lol my pleasure